OpenStudy (anonymous):

Solve for x, y, and z 5x +2y - 5z = -45 x+y + 5z = -25 x-7y-6z = 6

OpenStudy (anonymous):

x = -2390/207, y = -50/207, z = -547/207

OpenStudy (anonymous):

excuse me but can u help me out

OpenStudy (anonymous):

Check out : "Solving 3 Equations in 3 Unknowns" at the following web site: http://www.youtube.com/watch?v=woqq3Sls1d8